OpenStudy (anonymous):

The sum of three consecutive integers is 111. If "n" represents the least of the three integers, which of the following equations represents the statement above? 3n= 111 3n+2=111 3n+4=111 3n+6= 111 3n+9= 111

OpenStudy (anonymous):

let the integer be x , x+1 , x+ 2 x + x+1 + x+ 2 = 111 3x + 3 = 111

OpenStudy (anonymous):

none of the above:) 3x+3=111 x+1=37 is the answer:)